### v2

@@ 3.3 Application: two agents, quadratic costs
We know from Theorem 3.2 that we can restrict attention to SGE contracts. And with only two agents, the set of SGE contracts can be parametrized by a single parameter $\lambda$, where $f_{i}(S)=0$ whenever $i\notin S$, $f_{i}(S)=1$ whenever $S=\{i\}$, and $f_{1}(\{1,2\})=1-f_{2}(\{1,2\})=\lambda$. The equilibrium conditions for this contract from Equation (2) are
$C_{1}p_{1}$ $=1-p_{2}+\lambda p_{2}$
$C_{2}p_{2}$ $=1-p_{1}+(1-\lambda)p_{1}.$
For each $\lambda$, this system of equations has a unique solution:
@@
As Theorem 3.3 demonstrates, agent $1$’s share $\lambda_{1}(w)$ is increasing in $w$. In terms of comparative statics, this indicates that a principal whose objective places more weight on agent $1$ will give a larger share of the reward to agent $1$. Perhaps more surprisingly, $\lambda_{1}(1)=\frac{1}{2}$ regardless of the $C_{i}$, and since $\lambda_{1}(w)$ is increasing, it follows that the optimal contract will split the reward equally between the agents if the principal values their success equally, and in general will give a larger share of the reward to the agent whose success the principal values more. The robustness of this structure of the optimal contract to heterogeneity in the agents’ costs is intriguing, and an interesting open question is whether this robustness holds more generally, and if so, under what conditions.
3.4 Proof of Theorem 3.2
In this subsection, we will discuss the ideas and techniques for proving Theorem 3.2. The result naturally splits into two directions, and we discuss each of these in turn.
First, we show that SGE contracts induce maximal equilibria. Notice that if $f$ is a SGE contract, then $\mathbb{E}\left[\sum_{i\in S}{f_{i}(S)}\,|\,S\neq\emptyset\right]=1$, and thus, it follows from Lemma 3.1 that for any $p\in E(f)$, $z(p)=1$. To prove that $p$ is maximal, we show that $z(x)$ is strictly increasing in $x_{i}$ for $x\geq p$, and thus, for any $q\in[0,1]^{n}$ which coordinate-wise dominates $p$, $z(q)>z(p)$. But since we know from Corollary 1 that for any $q\in\mathcal{E}$, $z(q)\leq 1$, it follows that there is no equilibrium $q$ that dominates $p$.
{lemrep}
If $f\in\mathcal{F}_{SGE}$, then $p\in E(f)$ is maximal.
@@
Suppose $p$ is not maximal. Then, there must be a $q\in\mathcal{E}$ that dominates $p$. Observe that since $p\in E(f)$ and $f\in\mathcal{F}_{SGE}$, we have from Lemma 3.1 that $z(p)=1$. Also, since $q\in\mathcal{E}$, we have from Corollary 1 that $z(q)\leq 1$. Now, consider the following:
$\dfrac{\partial z}{\partial x_{i}}\Bigr{|}_{x\geq p}$ $=x_{i}c_{i}^{\prime\prime}(x_{i})+c_{i}^{\prime}(x_{i})-\prod_{j\neq i}(1-x_{j})$
$>c_{i}^{\prime}(x_{i})-\prod_{j\neq i}(1-x_{j})$ (because $(\text{because }c_{i}\text{ is convex})$
$\geq c_{i}^{\prime}(p_{i})-\prod_{j\neq i}(1-x_{j})$ (because $(\text{because }c_{i}\text{ is convex})$
$\geq c_{i}^{\prime}(p_{i})-\prod_{j\neq i}(1-p_{j})$ (because $(\text{because }x_{j}\geq p_{j})$
$\geq 0$ (because $(\text{because }p\in E(f))$
Note that the last inequality holds because we know from the first order condition in Equation (2) that $c_{i}^{\prime}(p_{i})=\mathbb{E}[f_{i}(S)|i\in S]\geq\prod_{j\neq i}(1-p_{j})$ since $f\in\mathcal{F}_{SGE}$. Now, since $q$ dominates $p$, $q\geq p$ and it follows from above that $z(q)>z(p)$. But this is a contradiction since $z(p)=1\geq z(p)$. It follows then that $p\in E(f)$ is maximal. ∎
Proving the other direction that maximal equilibria can only be induced by SGE contracts is quite involved and nontrivial in comparison to the single-agent case. To see why, consider a contract $f$ which is not SGE. We want to show that $p\in E(f)$ cannot be maximal. Intuitively, one might suspect that the principal can adjust the contract in such a way that $r_{i}(f,p_{-i})$ increases for some agent, either by decreasing its reward when it fails or by increasing its reward when it succeeds, and that this transformation should result in an equilibrium that dominates $p$. While this is true when there is only a single agent, the equilibrium effects with multiple agents mean that this transformation may simultaneously result in a decrease in the success probability chosen by another agent.
For an illustration, let us go back to the two agent example from Subsection 3.3. Consider the contract $f$ such that
$f_{i}(S)=\begin{cases}\beta_{i},&\text{ if }S=\{i\}\\ 0,&\text{ otherwise }\end{cases}$
That is, $f$ awards $\beta_{i}\in[0,1]$ to agent $i$ if and only if $i$ is the only successful agent. Using the first order conditions in Equation (2), one can verify that $f$ has a unique equilibrium $p=(p_{1},p_{2})$ with
$p_{i}=\dfrac{\beta_{i}(C_{-i}-\beta_{-i})}{C_{1}C_{2}-\beta_{1}\beta_{2}}.$
Now it is easy to check that increasing $\beta_{i}$ would increase $p_{i}$, but it may also be accompanied by a decrease in $p_{-i}$. Thus, given a contract that is not SGE, simply increasing the reward of an agent when it succeeds does not necessarily lead to coordinate-wise dominant equilibrium.
To prove that non-SGE contracts lead to equilibria that are not maximal, we essentially demonstrate that there is always some adjustment that the principal can make that leads to a dominant equilibrium. More precisely, given a non-SGE contract $f$ and $p\in E(f)$, we will illustrate transformations of $f$ that work when $f$ satisfies certain properties. And in case $f$ does not satisfy those properties, we will find a non-SGE contract $g\in E^{-1}(p)$ that does satisfy them, thus proving the result.
First, we show that in any maximal equilibrium, every agent chooses a strictly positive probability of success.
{lemrep}
If $p\in E(f)$ is such that $p_{k}=0$ for some $k\in[n]$, then $p$ is not maximal.
Proof.
Let $f\in E^{-1}(p)$ be an FGN contract, and suppose $p_{k}=0$. Consider the contract $g$ where $g_{i}(S)=f_{i}(S\setminus\{k\})$ for $i\neq k$ and
Let $p_{k}^{\prime}=b_{k}(g,p_{-i})$, and let $p^{\prime}=(p_{k}^{\prime},p_{-k})$. Observe that since $b_{k}(g,p_{-i})>0$, $p_{k}^{\prime}>0$, and for $i\neq k$, $r_{i}(g,p_{-i}^{\prime})=r_{i}(f,p_{-i})$, so $b_{i}(g,p_{-i^{\prime}})=b_{i}(f,p_{-i})=p_{i}$. Hence, $p^{\prime}\in E(g)$ and $p^{\prime}$ dominates $p$. Thus, if $p_{k}=0$ for some $k$, then $p\notin\mathcal{P}$, and the result follows. ∎
We prove this by constructing a new contract $g$ where $g_{k}(S)=\mathbbm{1}_{S=\{k\}}$ and $g_{i}(S)=f_{i}(S\setminus\{k\})$ for $i\neq k$. For agents other than $k$, this contract induces a game that is strategically equivalent to the game induced by $f$ when agent $k$ chooses $p_{k}=0$. In particular, switching from the contract $f$ to the contract $g$ induces agent $k$ to choose a strictly positive probability of success without inducing any change in the other agents’ behavior.
A useful consequence of Lemma 3.4 is that if $p$ is maximal, it must be that $10<p_{i}<$ 1 for all $i\in[n]$, and therefore, for any $S\subset[n]$, there is positive probability that $S$ is exactly the set of successful agents.
Next, we show that if $f\in\mathcal{F}_{FGN}$ never exhausts the entire budget, $p\in E(f)$ cannot be maximal. {lemrep} If $f\in\mathcal{F}_{FGN}$ is such that $\sum_{i\in S}f_{i}(S)<1$ for all $S\subset[n]$, then $p\in E(f)$ is not maximal.
Proof.
If $p_{i}=0$ for some $i\in[n]$, it follows from Lemma 3.4 that $p$ is not maximal. So suppose $p_{i}>0$ for all $i\in[n]$. Now let $p^{\prime}=(p_{1}+\epsilon,p_{2}+\epsilon,\dots,p_{n}+\epsilon)$. For each $i\in[n]$, let $t_{i}(\epsilon)$ solve
$c_{i}^{\prime}(p^{\prime}_{i})=t_{i}(\epsilon)\sum_{S\subset[n]_{-i}}(f_{i}(S\cup\{i\})){\Pr}^{[n]_{-i}}_{p^{\prime}_{-i}}(S).$
Observe that as $\epsilon\to 0$, $t_{i}(\epsilon)\to 1$ for all $i\in[n]$. Since $\sum_{i\in S}{f_{i}(S)}<1$ for all $S\subset[n]$ and $t_{i}(\epsilon)$ is continuous in $\epsilon$, we can find $\epsilon>0$ small enough so that the contract $f^{\prime}_{i}(S)=f_{i}(S)*t_{i}(\epsilon)$ for all $S\subset[n]$ and $i\in S$ is a feasible contract. By the definition of $t_{i}$, $p^{\prime}$ with $p_{i}^{\prime}=p_{i}+\epsilon$ for $i\in[n]$ will be an equilibrium under $f^{\prime}$. Thus, we have that $p$ is not maximal.
To prove this, for any $\varepsilon>0$, let $p^{\prime}$ be the profile with $p_{i}^{\prime}=p_{i}+\varepsilon$ for all $i$, let $t_{i}(\varepsilon)=\frac{c_{i}^{\prime}(p_{i}^{\prime})}{\mathbb{E}_{p^{\prime}}[f_{i}(S)\,|\,i\in S]}$, and let $g_{i}(S)=t_{i}(\varepsilon)\cdot f_{i}(S)$. Observe that if $\sum_{i\in S}g_{i}(S)\leq 1$ for all $S\subseteq[n]$, then $g$ is a contract with equilibrium $p^{\prime}$ which dominates $p$. Since $\lim_{\varepsilon\to 0}t_{i}(\varepsilon)=1$, $g$ is a contract for all sufficiently small $\varepsilon$, and hence $p$ is dominated.
Now, since Equation (2) is linear in $f$ and $\mathcal{F}_{FGN}$ is closed and convex, the set of FGN contracts that implement $p$ is closed and convex. This has the following immediate corollary: if for each nonempty $S\subseteq[n]$ there is an FGN contract $g^{S}$ which implements $p$ such that $\sum_{i\in S}g_{i}^{S}(S)<1$, then there is an FGN contract $f$ as in Lemma 3.4 that implements $p$. Moreover, a simple argument shows that if $p_{i}>0$ for all $i$ and there is some $j$ such that $\sum_{i\in S}f_{i}(S)<1$ for some $S\ni j$, then there is some FGN $g$ that implements $p$ such that $\sum_{i\in S}g_{i}(S)<1$ for all $S\ni j$. Putting these observations together provides a larger class of FGN contracts which cannot be optimal.
Corollary 2.
If $f\in\mathcal{F}_{FGN}$ is such that for each $j$ there is some $S\ni j$ such that $\sum_{i\in S}f_{i}(S)<1$, then $p\in E(f)$ is not maximal.
To prove Theorem 3.2, it remains to deal with contracts that are not SGE but do not satisfy the assumption of Corollary 2. This final step is involved, so we provide here only a high-level overview. Given an equilibrium $p$, consider the set of agents $i$ for whom $\sum_{j\in S}f_{j}(S)=1$ for every $S\ni i$ and every FGN contract $f$ that implements $p$. Starting with an FGN contract for $p$, we can then construct a new contract for $p$ such that the payoffs of these agents does not depend on the payoffs of the other agents. From this point, we can use the same techniques as above to modify the contract to induce the other agents to choose higher success probabilities. Hence, if $f$ is not an SGE contract, then any equilibrium for $f$ cannot be maximal.
{lemrep}
If $f\notin\mathcal{F}_{SGE}$, then $p\in E(f)$ is not maximal.
Proof.
Suppose $p\in\mathcal{P}$. Let
and let
$\kappa_{p}:=\{i\in[n]\,:\,\{i\}\in K_{p}\}.$
We now show that $K_{p}=\{\emptyset\}$.
  1. 1.
    Step 1: Suppose $S\in K_{p}$. For any $T\subset S$, $T\in K_{p}$.
    Let $f\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ be such that $\sum_{i\in S}{f_{i}(S)}<1$. If $\sum_{i\in T}{f_{i}(T)}<1$, we are done. Otherwise, pick an agent $i\in T$ such that $f_{i}(T)>0$ and consider a contract $g$ which differs from $f$ only in its award for agent $i$ at $S$ and $T$. In particular, let $g$ be such that $g_{i}(S)=f_{i}(S)+\epsilon$ and $g_{i}(T)=f_{i}(T)-\delta$ where $\epsilon,\delta>0$ are chosen so that $p\in E(g)$. Note that we can do this because we know from Lemma 3.4 that for all $i\in[n]$, $10<p_{i}<$ 1 and therefore, ${\Pr}^{[n]_{-i}}_{p_{-i}}(S)>0$ for all $i\in[n]$ and all $S\subset[n]_{-i}$. It follows then that $g\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ and $\sum_{i\in T}{g_{i}(T)}<1$. Thus, $T\in K_{p}$.
  2. 2.
    Step 2: Suppose $S,T\in K_{p}$. Then, $S\cup T\in K_{p}$.
    Let $f,g\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ be such that $\sum_{i\in S}{f_{i}(S)}<1$ and $\sum_{i\in T}{g_{i}(T)}<1$. Consider the contract $h=\frac{1}{2}f+\frac{1}{2}g$. Since Equation (2) is linear in $f$ and $\mathcal{F}_{FGN}$ is closed and convex, the set of FGN contracts that implement $p$ is closed and convex. Thus, $h\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ and also $\sum_{i\in S}{h_{i}(S)}<1$ and $\sum_{i\in T}{h_{i}(T)}<1$. Now, if $\sum_{i\in S\cup T}{h_{i}(S\cup T)}<1$, we are done. Otherwise, pick any agent $i\in S\cup T$ (WLOG, let $i\in S$) such that $h_{i}(S\cup T)>0$ and consider a contract $h^{\prime}$ which differs from $h$ only in its award for agent $i$ at $S\cup T$ and $S$. In particular, let $h^{\prime}$ be such that $h^{\prime}_{i}(S\cup T)=h_{i}(S\cup T)-\epsilon$ and $h^{\prime}_{i}(S)=h_{i}(S)+\delta$ where $\epsilon,\delta>0$ are chosen so that $p\in E(h^{\prime})$. Again, we can do this because we know that for all $i\in[n]$, $10<p_{i}<$ 1. It follows then that $h^{\prime}\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ and $\sum_{i\in S\cup T}{h^{\prime}_{i}(S\cup T)}<1$. Thus, $S\cup T\in K_{p}$.
    Note that it follows from Steps 1 and 2 that $K_{p}=2^{\kappa_{p}}$.
  3. 3.
    Suppose towards a contradiction that there is an $S\subset[n]$ such that $\kappa_{p}^{C}\cap S\neq\phi$ and $f_{i}(S)>0$ for some $i\in\kappa_{p}$. Let $g\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ be such that $g_{i}(\{i\})<1$. Consider the contract $h=\frac{1}{2}f+\frac{1}{2}g$. Then $h\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ and also $h_{i}(S)>0$ and $h_{i}(\{i\})<1$. Now, consider a contract $h^{\prime}$ which differs from $h$ only in its award for agent $i$ at $S$ and $\{i\}$. In particular, let $h^{\prime}$ be such that $h^{\prime}_{i}(\{i\})=h_{i}(\{i\})+\epsilon$ and $h^{\prime}_{i}(S)=h_{i}(S)-\delta$ where $\epsilon,\delta>0$ are chosen so that $p\in E(h^{\prime})$. We can do this because for all $i\in[n]$, $10<p_{i}<$ 1. It follows then that $h^{\prime}\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ and $\sum_{i\in S}{h^{\prime}_{i}(S)}<1$. But this means that $S\subset\kappa_{p}$ which is a contradiction.
  4. 4.
    Step 4: Suppose $\kappa_{p}\neq\phi$. Then there is a $p^{\prime}$ that Pareto dominates $p$.
    by averaging over the outcomes of agents in $\kappa_{p}$ under $g$.
    $c_{i}^{\prime}(p^{\prime}_{i})=t_{i}(\epsilon)\sum_{S\subset A_{-i}}(h_{i}(S\cup\{i\})){\Pr}^{A_{-i}}_{p^{\prime}_{-i}}(S)$
    Observe that as $\epsilon\to 0$, $t_{i}(\epsilon)\to 1$ for all $i\in A$. Since $\sum_{i\in S}{h_{i}(S)}<1$ for all $S\subset A$ and $t_{i}(\epsilon)$ is continuous in $\epsilon$, we can find $\epsilon>0$ small enough so that the contract $h^{\prime}_{i}(S)=h_{i}(S)*t_{i}(\epsilon)$ for all $S\subset A$ and $i\in S$ is a feasible contract. By the definition of $t_{i}$, $p^{\prime}$ with $p_{i}^{\prime}=p_{i}+\epsilon$ for $i\in\kappa_{p}$ and $p_{i}^{\prime}=p_{i}$ for $i\notin\kappa_{p}$ will be an equilibrium under $h^{\prime}$. Thus, we have that $p$ is not Pareto optimal.
It follows then that $K_{p}=\{\emptyset\}$. By definition of $K_{p}$, this means that for any $f\in E^{-1}(p)$, either $f\notin\mathcal{F}_{FGN}$ or $f\in\mathcal{F}_{SGE}$. Now suppose there exists an $f$ such that $f\notin\mathcal{F}_{FGN}$ and $p\in E(f)$. From Proposition 3.1, we can find $g\in\mathcal{F}_{FGN}$ such that $p\in E(g)$. Moreover, we know from the construction of $g$ in the argument of Proposition 3.1 that $g\notin\mathcal{F}_{SGE}$. Thus, we have that $g\in\mathcal{F}_{FGN},g\notin\mathcal{F}_{SGE},$ and $p\in E(g)$. But this means that $K_{p}\neq\emptyset$ which is a contradiction. Therefore, it must be that for any $f\in E^{-1}(p)$, $f\in\mathcal{F}_{SGE}$.
4 Implementing the maximal set
As Theorem 3.2 shows, maximal equilibria are characterized by the equilibria of SGE contracts, and so it is sufficient for the principal to optimize over SGE contracts. But observe that the set of SGE contracts is $\Theta(n2^{n})$-dimensional. Given the high dimensionality of the set of SGE contracts, optimizing over SGE contracts may still present the principal with a computationally difficult problem, and the optimal contract may be difficult to understand and implement. In this section, we identify a subclass of SGE contracts that is actually sufficient to implement the maximal set.
4.1 Luce contracts
We begin by introducing a class of weighted contracts, which provide a natural generalization of the equal-split-amongst-successful contract by allowing for weighted splits.
Definition 4.
Weighted contracts assigns weights to the agents and reward each successful agent with a fraction of the budget proportional to her weight. They are, however, insufficient to implement the maximal set. To see why, recall the structure of the optimal contract for the two agent problem from Theorem 3.3. For the objective $V(p)=wp_{1}+p_{2}$ with $w$ large, the optimal contract was such that if agent $1$ succeeded, it was awarded the entire budget regardless of whether agent $2$ succeeded or failed. And only in case agent $1$ failed, agent $2$ was awarded the entire budget if it was successful. It is easy to check that this optimal contract, while clearly implementing something in the maximal set, is not a weighted contract.
While this contract is not itself a weighted contract, it is a limit of weighted contracts: any sequence of weighted contracts with weights $(\lambda_{1},\lambda_{2})$ such that $\frac{\lambda_{1}}{\lambda_{2}}\to\infty$ approaches this contract in $\mathcal{F}$. Observe that this contract can be naturally interpreted as assigning a priority order over the agents and rewarding the highest-priority successful agent with the entire budget. In the general context with $n$ agents, any sequence of weighted contracts which converges must converge to a priority-based weighted contract, and we refer to such contracts as Luce contracts.
Definition 5.
[…]
Luce contracts reward each highest-priority successful agent with a fraction of the budget proportional to her weight. The choice of terminology is motivated by the analogy between the structure of Luce contracts and the selection rules that satisfy Luce’s choice axiom (Luce [30]). Just as Luce rules have the property that the relative odds of selecting one item over another from a pool is unaffected by which other items are in the pool, Luce contracts have the property that the reward of a successful agent relative to another successful agent is independent of which other agents succeed. While weighted contracts alone are insufficient, the following result shows that Luce contracts are indeed sufficient to implement the maximal set. {thm} If $p\in\mathcal{P}$, there is a unique Luce contract $f\in\mathcal{F}_{Luce}$ such that $p\in E(f)$.
In words, if $p$ is maximal, then there is a unique Luce contract that implements $p$. Note that since this contract is unique, the set of Luce contracts is minimal among sets of contracts that implement the maximal set. As a consequence of Theorem 5, it follows that it is sufficient for the principal to optimize over the set of Luce contracts.
Corollary 3.
For any strictly increasing, continuous objective $V(p)$,
4.2 Proof of Theorem 5
To prove Theorem 5, we first characterize all profiles $p\in(0,1)^{n}$ that can be implemented by a Luce contract $f\in\mathcal{F}_{Luce}$ with some budget.
To begin, suppose $p\in(0,1)^{n}$ is an equilibrium of a Luce contract $f$ with budget $B$. Since $f\in\mathcal{F}_{FGN}$, we have from Equation (3) that the expected total payment from principal to agents in $I\subset[n]$ is
$\mathbb{E}\left[\sum_{i\in I}B\cdot f_{i}(S)\right]=\sum_{i\in I}p_{i}\cdot c_{i}^{\prime}(p_{i}).$
And since $f\in\mathcal{F}_{Luce}$, we have by an alternative calculation that
$\mathbb{E}\left[\sum_{i\in I}B\cdot f_{i}(S)\right]\leq B\cdot\mathbb{P}[S\cap I\neq\emptyset],$
and in particular, for $I=[n]$,
$\mathbb{E}\left[\sum_{i\in[n]}B\cdot f_{i}(S)\right]=B\cdot\mathbb{P}[S\neq\emptyset].$
Hence, for any subset $I$ of the agents,
$\dfrac{\sum_{i\in I}p_{i}\cdot c_{i}^{\prime}(p_{i})}{\sum_{i\in[n]}p_{i}\cdot c_{i}^{\prime}(p_{i})}\,\,\leq\,\,\dfrac{\mathbb{P}[S\cap I\neq\emptyset]}{\mathbb{P}[S\neq\emptyset]}\cdot$ (5)
This provides necessary conditions for any $p\in(0,1)^{n}$ to be implementable by a Luce contract. The following result shows that this is also sufficient: any $p\in(0,1)^{n}$ which satisfies (5) can be implemented by a Luce contract.
{proprep}
Suppose $p\in(0,1)^{n}$. There exists a Luce contract $f\in\mathcal{F}_{Luce}$ that implements $p$ if and only if for all $I\subset[n]$,
$\dfrac{\sum_{i\in I}p_{i}\cdot c_{i}^{\prime}(p_{i})}{\sum_{i\in[n]}p_{i}\cdot c_{i}^{\prime}(p_{i})}\,\,\leq\,\,\dfrac{\mathbb{P}[S\cap I\neq\emptyset]}{\mathbb{P}[S\neq\emptyset]}\cdot$
Proof.
Assume there is a Luce contract $f\in\mathcal{F}_{Luce}$ with budget $B>0$ that implements $p$. Then for any $I\subseteq[n]$, the expected total reward to agents in $I$ is
$\sum_{i\in I}p_{i}\cdot c_{i}^{\prime}(p_{i})=\mathbb{E}\left[\sum_{i\in I}B\cdot f_{i}(S)\right]=B\cdot\mathbb{E}\left[\sum_{i\in I}f_{i}(S)\,|\,S\cap I\neq\emptyset\right]\cdot\mathbb{P}[S\cap I\neq\emptyset],$
and since $f$ is a Luce contract with budget $B$,
$\mathbb{E}\left[\sum_{i\in I}f_{i}(S)\,|\,S\cap I\neq\emptyset\right]\leq 1$
with equality if $I=[n]$. Hence, if there is a Luce contract with budget $B$ that implements $p$, then (5) must hold.
Now, suppose that $p\in(0,1)^{n}$ satisfies (5). Observe that for any Luce contract that uses a budget of $B=\dfrac{\sum_{i\in[n]}p_{i}\cdot c_{i}^{\prime}(p_{i})}{\mathbb{P}[S\neq\emptyset]}$,
$\sum_{j\in[n]}p_{j}\cdot c_{j}^{\prime}\left(b_{j}\left(f,p_{-j}\right)\right)=B\cdot\mathbb{E}[\sum_{j\in[n]}f_{j}(S)]=B\cdot\mathbb{P}[S\neq\emptyset]=\sum_{j\in[n]}{p_{j}\cdot c_{j}^{\prime}(p_{j})},$
so
Moreover, for any agent $i$,
$\sum_{j:j\succ i}p_{j}\cdot c_{j}^{\prime}\left(b_{j}\left(f,p_{-j}\right)\right)=B\cdot\mathbb{E}[\sum_{j:j\succ i}f_{j}(S)]=B\cdot\mathbb{P}[S\cap\{j:j\succ i\}\neq\emptyset]\geq\sum_{j:j\succ i}p_{j}\cdot c_{j}^{\prime}(p_{j}),$
so
Let
@@ 4.2 Proof of Theorem 5
$\mathcal{C}(f):=\{i\in[n]\,:\,c_{i}^{\prime}(b_{i}(f,p_{-i}))-c_{i}^{\prime}(p_{i})=Z(f)\}$
From above, we know that $Z(f)\geq 0$ for all $f\in\mathcal{F}_{Luce}$ and so
@@
which implies $f\in E^{-1}(p)$. Thus, for any $p$ that satisfies (5), there is a Luce contract $f$ and a budget $B$ such that $p\in E(f)$.
In particular, for any $p\in(0,1)^{n}$ that satisfies (5), the principal can induce $p$ by setting aside a budget of
$B=\dfrac{\sum_{i\in[n]}p_{i}\cdot c_{i}^{\prime}(p_{i})}{\mathbb{P}[S\neq\emptyset]},$
and committing to distributing this budget as per an appropriately defined Luce contract. The proof of Proposition 4.2 is technical and relegated to the appendix.
Now we turn to the proof of Theorem 5. The result states that under the small budget assumption (normalized to $B=1$), if $p$ is maximal, then there is a unique Luce contract that implements $p$. From Theorem 3.2, we know that $p$ must be an equilibrium of a SGE contract. But then, by the same calculation as above, it follows that $p$ must satisfy (5). From Proposition 4.2, there is a Luce contract that implements $p$. Thus, we get that every maximal equilibrium can be implemented by a Luce contract.
We now show that there is a unique Luce contract that implements $p$. To this end, suppose that $f$ and $g$ are distinct weighted contracts which both implement $p$, and let $j$ be an agent whose ratio of weights $\frac{\lambda_{j}^{f}}{\lambda_{j}^{g}}$ is minimal. Then $f_{j}(S)\leq g_{j}(S)$ for all $S\subseteq[n]$ with strict inequality for some $S$, so the best response of agent $j$ to $p_{-j}$ must be strictly lower under $f$ than under $g$, and hence $p$ cannot be an equilibrium for both $f$ and $g$. Though more technical, a similar argument extends to Luce contracts.
{lemrep}
Suppose $f\in\mathcal{F}_{Luce}$ and $p\in E(f)$. Then for any $g\in\mathcal{F}_{Luce}$ such that $g\neq f$, $p\notin E(g)$.

### v3

@@ 3 Main results
We know from Theorem 3 that we can restrict attention to SGE contracts. And with only two agents, the set of SGE contracts can be parametrized by a single parameter $\lambda$, where $f_{i}(S)=0$ whenever $i\notin S$, $f_{i}(S)=1$ whenever $S=\{i\}$, and $f_{1}(\{1,2\})=1-f_{2}(\{1,2\})=\lambda$. The equilibrium conditions for this contract from Equation (3) are
$C_{1}p_{1}$ $=(1-p_{2}+\lambda p_{2})\cdot B$
$C_{2}p_{2}$ $=(1-p_{1}+(1-\lambda)p_{1})\cdot B.$
Going forward, we normalize $C_{i}=\frac{C_{i}}{B}$. For each $\lambda$, this system of equations has a unique solution:
@@
Proposition 3 shows that an agent’s share of the budget is increasing in the relative weight that the principal places on that agent’s success. Perhaps more surprisingly, the result also prescribes an equal distribution of the budget whenever the principal values agents’ successes equally ($\lambda_{1}(1)=\frac{1}{2}$), regardless of heterogeneity in their cost parameters $C_{1},C_{2}$. Moreover, because $\lambda_{1}(w)$ is increasing, the optimal contract always awards a greater share of the reward to the agent whose success the principal values more, independently of the agents’ costs. This robustness of the structure of optimal contract to cost heterogeneity is intriguing. An interesting open question is whether this robustness property holds more broadly, and if so, under what general conditions.
4 Luce contracts without budget constraints
In the previous section, we established that Luce contracts are optimal in environments where the principal has access to an exclusive-use budget. However, this assumption departs from standard contract-design models, where rewarding agents entails a direct cost, and the principal must weigh this cost against the benefits of inducing higher effort. In this section, we argue that Luce contracts remain desirable even when the principal is not budget-constrained.
In the absence of a budget constraint, the principal has access to a broad range of contracts for implementing any desired effort profile $q\in(0,1)^{n}$. For instance, the principal could offer each agent $i\in[n]$ an independent piece-rate contract that pays $c_{i}^{\prime}(q_{i})$ upon success and nothing otherwise. The principal could also offer a bonus-pool contract which pays $\dfrac{q_{i}c_{i}^{\prime}(q_{i})}{\Pi_{i\in[n]}q_{i}}$ to agent $i\in[n]$ if all agents succeed, and nothing otherwise. Perhaps, the principal could also just set aside a budget $B=\frac{\sum_{i\in[n]}q_{i}\cdot c_{i}^{\prime}(q_{i})}{\mathbb{P}_{q}[S\neq\emptyset]}$, and design a Luce contract that implements $q$. Unfortunately, this may not always be feasible, as the structure of Luce contracts introduces constraints on effort profiles that they can implement. The following result characterizes effort profiles for which implementation via a Luce contract is feasible.
{proprep}
Suppose $p\in(0,1)^{n}$. There exists a Luce contract $f\in\mathcal{F}_{Luce}$ with budget $B>0$ that implements $p$ if and only if for all $I\subset[n]$,
$\sum_{i\in I}p_{i}\cdot c_{i}^{\prime}(p_{i})\leq\mathbb{P}_{p}[S\cap I\neq\emptyset]\cdot B,$ (5)
where $B=\dfrac{\sum_{i\in[n]}p_{i}\cdot c_{i}^{\prime}(p_{i})}{\mathbb{P}_{p}[S\neq\emptyset]}$.
@@ 4 Luce contracts without budget constraints
Suppose $f\in\mathcal{F}_{Luce}$ with budget $B>0$ implements $p$. Then for any $I\subseteq[n]$, the expected total reward to agents in $I$ is
$\sum_{i\in I}p_{i}\cdot c_{i}^{\prime}(p_{i})$ $=\mathbb{E}_{p}\left[\sum_{i\in I}f_{i}(S)\right]\cdot B$
$\leq\mathbb{P}_{p}[S\cap I\neq\emptyset]\cdot B$
with equality for $I=[n]$. Thus, $p$ must satisfy (5).
Now, suppose $p\in(0,1)^{n}$ satisfies (5). For any $f\in\mathcal{F}_{Luce}$ with budget $B$, let
@@
$\mathcal{C}(f):=\{i\in[n]\,:\,c_{i}^{\prime}(b_{i}(f,p_{-i}))-c_{i}^{\prime}(p_{i})=Z(f)\}.$
From Equation (3),
Moreover, for any agent $i\in[n]$,
It follows that $Z(f)\geq 0$ for all $f\in\mathcal{F}_{Luce}$ and so
@@
which implies $f\in E^{-1}(p)$. Thus, for any $p$ that satisfies (5), there is a Luce contract $f$ with budget $B$ such that $p\in E(f)$.
When possible, Luce contracts provide a particularly desirable alternative for implementation over other contracts by mitigating the uncertainty in total payments. For a contract $f$ with budget $B$ and a profile $p$, denote by $T_{f,B}=\sum_{i}f_{i}(S)\cdot B$ the (random) total payment.
{thmrep}
Suppose $p\in(0,1)^{n}$ is implemented by a Luce contract $f$ with budget $B$ and an FGN contract $g$ with budget $B^{\prime}$. Then $T_{g,B^{\prime}}$ is a mean preserving spread of $T_{f,B}$ at $p$.